# Geometric Algebra Meets Large Language Models: Instruction-Based Transformations of Separate Meshes in 3D, Interactive and Controllable Scenes

### v1

@@ 4.2. Collision Detection Module
Similar to Yang et al. (2023), our system conducts a Depth-First Search (DFS) on a constructed 3D grid surrounding the colliding objects, identifying the first valid solution within the considered search space. We only consider collisions that can be determined using the bounding box information for each object. To prevent placing objects directly on the corners of the 3D grid, we incorporate a fixed buffer zone to ensure adequate spacing during object placement. Each object is described by 6 parameters, corresponding to its axis-aligned bounding box : $(x,y,z)$ for the center coordinates, $w,d$ and $h$ for the width, depth and height. The process starts with initial placements of the target objects and concludes at the first occurrence of a valid placement. We prioritize solutions that are close to the original configurations, exploring all possibilities and gradually allowing for more distant solutions in a linear progression. Our goal is to minimize disruptions from the initial configuration while preserving the spatial integrity of the scene. This collision resolution operates within a set timeframe (e.g., 0.5 or 1 second). Finally, all transformations are applied.
5. Evaluation
@@
We conduct comprehensive human evaluations to assess shenlong in object repositioning, engaging 20 annotators: 5 3D designers, 5 game programmers, and 10 individuals outside the gaming and 3D design discipline areas. This study examines shenlong’s proficiency in editing diverse scenes. Through these user studies, we demonstrate that, by harnessing CGA in our prompting strategies, we significantly outperform baseline alternatives, including those made with NVIDIA Omniverse (see Section 5.2.1).
5.1. Experimental Setup
@@
We generated ten diverse scenes using the HOLODECK framework (Yang et al., 2023) for human evaluation, encompassing various settings including living rooms, wine cellars, kitchens, and medical operating rooms. To facilitate the integration of diverse scenes into the TDW framework, we developed a custom importer. For each scene, we created five variations of templated queries, populating them with objects randomly selected from within the scene. We evaluated 50 distinct queries for each scene, resulting in a total of 2,500 scenes (coupled with prompts) for human assessment. For each query, we compared the initial scene with the resulting scene after processing by the two baseline system and shenlong. Examples of the human evaluation are presented in the Supplementary Material. For each edited scene, we displayed top-down view images and a 360-degree video view, asking annotators to assess the accuracy of the editing performed. Each object repositioning query was evaluated by five annotators, and a result was considered valid only if all annotators unanimously agreed on the assessment. Due to space limitations, the evaluation queries only are displayed on the y-axis of Figure 4. It should be noted that all preliminary experiments on dilation-related queries with shenlong and other baselines achieved a perfect success rate, leading us to omit these queries. The queries are categorized into five groups, each containing ten queries that evaluate the system’s performance across a progressively increasing difficulty gradient. Below, we provide further details on the group categories used in our analysis:
@@
The classification of “simple” and “composite” queries is based on the study of Manesh et al. (2024) on natural language in virtual environment creation. Building on this, our work further explores fuzzy and more complex scenarios involving the repositioning of multiple objects.
5.1.1. System Configuration
@@ 5.2. Benchmarking
5.2.1. Baseline Systems
One of our baseline systems, for comparison is one of the publicly available prompts from NVIDIA’s Omniverse [^1] https://github.com/NVIDIA-Omniverse/kit-extension-sample-airoomgenerator/blob/main/exts/omni.example.airoomgenerator/omni/example/airoomgenerator/prompts.py. This prompt operates by generating a JSON output that details object placements within a 3D space, similar to our system. Contrary to shenlong, the Omniverse prompt imposes no constraints on the reasoning scheme that the LLM should follow, requesting only the final positions of the objects. The Omniverse prompt accepts input specifications for each object, including its name, dimensions along the $X$, $Y$ and $Z$ axes, and a centrally located origin point. To ensure this baseline prompt is comparable to our system, we include only points of interest relevant to the scene editing query, rather than the entire scene description as initially done. This refined input approach significantly impacts the LLM’s response time. Our experiments demonstrate that providing only the relevant points of interest leads to a substantial decrease in response time—specifically, an average of $3.3\pm 0.1$ times faster—without compromising overall accuracy. Due to these findings, we report results exclusively from this optimized methodology. Finally, we have extended the prompt by incorporating Euler angles for object orientation.
Figure 3. The left chart compares the average success rates, and the right chart compares the average LLM response times for three different systems—Omniverse, Euclidean, and CGA—across five categories of queries: Simple Queries, Compositional Queries, Fuzzy Queries, Compositional Fuzzy Queries, and Hard Queries. Success rates are measured on a scale from 0.0 to 1.0. Horizontal dashed lines indicate the overall mean success rate and response times for each system.
Our second baseline model is derived from recent work showing that LLMs can predict a dense sequence of end-effector poses for manipulation tasks (Kwon et al., 2024). We extend this concept to object repositioning on 3D scenes, treating it as a form of zero-shot trajectory generation. In this context, we task the LLM with constructing the necessary translation and rotation matrices to define the final trajectory, adapting the underlying techniques to suit scene manipulation objectives. We consider this baseline model to be the closest to our approach, as it guides the LLM to operate using templated functions that correspond to translation and rotation functions within the Euclidean space. This method aligns with our use of structured prompts that direct the LLM to generate specific geometric transformations necessary for scene editing tasks.
@@ 5.2.2. Performance Metrics
$S=\frac{1}{M}\sum_{i=1}^{M}\chi_{\text{correct}}(i)$
@@ 5.2.3. Results & Discussion
Figure 4. Query resolution performance for different prompts—Omniverse, Euclidean, and CGA—across five query categories: Simple, Compositional, Fuzzy, Compositional Fuzzy, and Hard. Success rate is assessed for each templated query over ten different scenes, with five query variations per scene.
@@
6. For coordinate extraction, employ the inner product operator |. For instance, use (X1.max | e2) * e2 to extract the Y-
axis coordinate.
7. CONSIDER the current position of objects when moving one in relation to another. Calculate transformations using the
initial locations derived from their bounding box points.
@@
3. **Input:** Translate X2 by 1 unit to the right and rotate it by 90 degrees around the z-axis.’
**Output:** {’X2’: T(1 * e1) * R(np.pi/2, e1, e2)’}
**Explanation:** The bounding box of X2 is defined by the points X2.max and X2.min. To translate X2 by 1 unit to the right, we must add 1 to the X-axis coordinate of both points. Therefore, the translation is T(1 * e1). The rotation rotor R(theta, u, v) rotates a vector by an angle theta in the plane defined by u ^ v. In this case, the rotation is
90 degrees in the z-axis is defined as the rotation around e1, e2 plane. Therefore, the rotation is R(np.pi/2, e1, e2). Since the rotation must be applied after the translation, the transformation is T(1 * e1) * R(np.pi/2, e1, e2).
4. **Input:** Move X1 on top of X0.’
**Output:** {’X1’:’T((X0.max+X0.min)/2 - (X1.max+X1.min)/2 + ((X0.max - X0.min) | e2) * e2)’}
**Explanation:** The bounding box of X1 is defined by the points X1.max and X1.min. The bounding box of X0 is defined by the points X0.max and X0.min. To move X1 on top of X0, we must translate X1 by the difference between the center of X0 and X1 and the Y-axis coordinate of X0. Therefore, the transformation is T((X0.max+X0.min)/2 - (X1.max+X1.min)/2 + ((X0.max - X0.min) | e2) * e2).
5. **Input:** Move X1 next to X0.’’
**Output:** {’X1’:’T((X0.max+X0.min)/2 - (X1.max+X1.min)/2 + (X0.max + (X0.max - X0.min) | e1) * e1)’}
**Explanation:** The bounding box of X1 is defined by the points X1.max and X1.min. The bounding box of X0 is defined by the points X0.max and X0.min. To move X1 next to X0, we must translate X1 by the difference between the center of
X0 and X1 and the X-axis coordinate of X0 and then add an offset along the e1 plane in order for the objects to not collide with each other. Therefore, the transformation is T((X0.max+X0.min)/2 - (X1.max+X1.min)/2 + ((X0.max - X0.min) | e1) * e1).
@@
- DOCUMENT Each Step: For every action you perform, you MUST provide a clear and
numbered list of operations used during your reasoning.
@@
- Its **IMPORTANT** to take into consideration all axis when performing relative transformations. Focus on the X **AND**
Z axis when moving objects around the room. Make changes in the Y axis when a vertical movement is required. So, for
example, when moving an object next to another, you should move it in both the X and Z axis, and when moving an
object on top of another, you should move it in the X and Z axis, and also in the Y axis.
- When no axis is defined for rotations, the default axis is the Y-axis which means a
rotation in the e1, e3 planes.
- When you want to move or rotate an object along an axis you must use the inner product operator |. For example, to move
an object along the Y-axis you must use (X1.max | e2) * e2.
- When no unit is specified for the translation, the default unit is 1 and for the rotation, the default unit is 90
degrees.
Figure 5. Template for Conformal Geometric Algebra Prompts Used by shenlong.
Performance Analysis by Query Group
Figure 3 presents a detailed evaluation of the prompts’ performance across different groups of queries. Each group’s results are analyzed to highlight specific strengths and weaknesses of the system in handling varying complexities. Specifically, the effectiveness and efficiency of the CGA, Omniverse, and Euclidean prompts were evaluated across the five categories of queries: Simple, Compositional, Fuzzy, Compositional Fuzzy, and Hard. We focused on two key performance metrics: average success rate and average response time, providing insights into each system’s scene editing capabilities.
Overall, CGA achieves the highest average success rate of 0.80 and the lowest average response time of 18 seconds (p ¡ 0.05), with the results being statistically significant. Interestingly, all prompts exhibited the same overall standard deviation, both in terms of success rate and response time, indicating consistent variability across the different editors. To assess the differences between the systems, we performed Student’s t-tests. In comparison, the Omniverse and Euclidean prompts achieve success rates of 0.74 and 0.72, and average response times of 21.5 seconds and 24.5 seconds, respectively. While the differences in success rates among Omniverse and Euclidean prompts are not statistically significant, the variations in average response times are. We conjecture that the lack of statistical significance in terms of success rates among the systems can be attributed to the nature of the reasoning methods employed. Both the Omniverse and Euclidean prompts rely on classical geometrical reasoning, which may lead to similar levels of performance. In contrast, shenlong utilizes a different approach to reasoning, which not only reflects in its superior performance but also in its statistical significance. This suggests that the distinct reasoning method employed by our system may contribute to its enhanced effectiveness.
@@
Detailed Performance by Individual Query
Figure 4 delves into the performance metrics for each individual query within the groups. This detailed breakdown provides insights into the prompts’ consistency, efficiency, and accuracy in executing scene editing tasks with varying degrees of complexity, facilitating a granular understanding of its operational effectiveness. With regard to the Simple Queries, all systems generally exhibit high success rates for straightforward tasks even in geometrically nuanced rotations such as those over specific planes, e.g., Rotate X1 over the e1 plane, with CGA consistently achieving perfect scores. This highlights CGA’s superior handling of precise geometric transformations. The Compositional Queries, which involve a combination of movements and rotations, we see CGA and Euclidean performing robustly, indicating effective integration of complex instructions. Omniverse, however, shows variability, particularly underperforming in scenarios like Rotate over the e3 plane and move X1 to the right where it achieves a 0.7 success rate compared to 1.0 by CGA.
In dealing with spatially ambiguous commands, i.e., Fuzzy Queries, Omniverse and CGA outperform Euclidean, especially notable in queries like Move X1 away from X2 where Euclidean drops to 0 against Omniverse’s perfect score. Interestingly, the Omniverse prompt struggles with the seemingly simple query, Move X1 on the left of X2, displaying a stark contrast to the near-perfect performance of the other systems. Although Omniverse is not presented with a specific move below example (only on top is provided in their prompts), all systems show generalization capabilities, with CGA achieving the highest success rate of 0.6. Similar generalization is observed with the Swap X1 and X2 query, where all systems are expected to perform well given their operational logic at the coordinate level; however, Omniverse shows surprisingly lower performance. In contrast, while the Euclidean and CGA systems perform less effectively on the Move X1 near X2 query, Omniverse excels with perfect success rates. Lastly, all prompts consistently fail the Rotate X1 to face X2 query. Intriguingly, despite being equipped with Euler angles, Omniverse also fails, suggesting a significant challenge area for current scene editing technologies.
The Compositional Fuzzy Queries category presents a notable challenge, combining fuzzy directives with multiple actions. In this category, Omniverse and CGA show comparable performances, indicating their effective handling of combined directives. However, it is important to note discrepancies in the Euclidean system’s performance. While it achieves good results on individual components of these queries, it performs poorly when actions are combined, as seen in queries like Move X1 next to X2 and rotate X3 by 90 degrees over the e2 plane and Swap X1 and X2 and move X3 on the left by 1 unit. This inconsistency highlights potential limitations in the Euclidean prompt’s ability to effectively integrate multiple spatial manipulations within a single query.
All systems exhibit challenges with the most demanding Hard Queries, which require intricate arrangements and precise manipulations. Notably, the Euclidean prompt displays extremely inconsistent behavior: it achieves perfect success rates on complex queries such as Place X1 on top of X2 and X3 below X2 and Align X1, X2 and X3 in a straight line, where other systems face difficulties. Conversely, it presents a zero success rate on tasks like Distribute X1, X2, X3 equally in a circle around X4 or Arrange X1, X2, X3 in a vertical line, then rotate them by 45 degrees keeping X1 intact, where other systems manage average performances. This variability suggests that while the Euclidean prompt excels in certain types of spatial reasoning, it may lack robustness in scenarios requiring dynamic spatial transformations or uniform object distribution. These findings underscore the need for further optimization and testing of the Euclidean system to enhance its consistency across a broader range of complex scene editing tasks. In summary, while CGA emerges as the most capable across a broad range of tasks, challenges remain, particularly in Hard and Compositional Fuzzy queries, underscoring the need for further advancements in LLM reasoning capabilities.
6. Limitations & Future Work
Object Representations
@@ 6. Limitations & Future Work
Currently, our system utilizes multiple intermediate representations of objects, including textual, templated, and bounding box references, requiring exact object names. To enhance this, we plan to implement semantic similarity measures at the token level or more advanced similarity searches using distributed representations. Additionally, we aim to enrich the system with further object information, such as orientation information, to better handle sophisticated queries involving relative rotations. By improving the system’s grasp of spatial relationships, improved performance in complex scenarios is expected.
System Optimisations
@@
Since we preprocess and create templated queries, it is straightforward to perform query caching optimizations (Mavrogiannis et al., 2023). We aim to implement query caching optimisations to enhance both the accuracy and responsiveness of our system. We also envision transforming the collision module into a multimodal agent (Yang et al., 2023), enabling nuanced handling of complex scenarios. The agent could be provided with a top-down view of the scene to facilitate finding better resolutions. Lastly, we plan to integrate speech interaction capabilities into our system.
Spatial Reasoning

### v2

@@ IV-B Collision Detection Module
Similar to [51], our system conducts a Depth-First Search (DFS) on a constructed 3D grid surrounding the colliding objects, identifying the first valid solution within the considered search space. We only consider collisions that can be determined using the bounding box information for each object. To prevent placing objects directly on the corners of the 3D grid, we incorporate a fixed buffer zone to ensure adequate spacing during object placement. Each object is described by 6 parameters, corresponding to its axis-aligned bounding box : $(x,y,z)$ for the center coordinates, $w,d$ and $h$ for the width, depth and height. The process starts with initial placements of the target objects and concludes at the first occurrence of a valid placement. We prioritize solutions that are close to the original configurations, exploring all possibilities and gradually allowing for more distant solutions in a linear progression. Our goal is to minimize disruptions from the initial configuration while preserving the spatial integrity of the scene. This collision resolution operates within a set timeframe (e.g., 0.5 or 1 second). Finally, all transformations are applied.
V Evaluation
@@
We conduct comprehensive human evaluations to assess Shenlong in object repositioning, engaging 20 annotators: 5 3D designers, 5 game programmers, and 10 individuals outside the gaming and 3D design discipline areas. This study examines Shenlong’s proficiency in editing diverse scenes. Through these user studies, we demonstrate that, by harnessing CGA in our prompting strategies, we significantly outperform baseline alternatives, including those made with NVIDIA Omniverse (see Section V-B1).
V-A Experimental Setup
@@
We generated ten diverse scenes using the HOLODECK framework [51] for human evaluation, encompassing various settings including living rooms, wine cellars, kitchens, and medical operating rooms. To facilitate the integration of diverse scenes into the TDW framework, we developed a custom importer. For each scene, we created five variations of templated queries, populating them with objects randomly selected from within the scene. We evaluated 50 distinct queries for each scene, resulting in a total of 2,500 scenes (coupled with prompts) for human assessment. For each query, we compared the initial scene with the resulting scene after processing by the two baseline system and Shenlong. Examples of the human evaluation are presented in the Supplementary Material. For each edited scene, we displayed top-down view images and a 360-degree video view, asking annotators to assess the accuracy of the editing performed. Each object repositioning query was evaluated by five annotators, and a result was considered valid only if all annotators unanimously agreed on the assessment. It should be noted that all preliminary experiments on dilation-related queries with Shenlong and other baselines achieved a perfect success rate, leading us to omit these queries. The queries are categorized into five groups, each containing ten queries that evaluate the system’s performance across a progressively increasing difficulty gradient. Below, we provide further details on the group categories used in our analysis:
@@
The classification of “simple” and “composite” queries is based on the study of [35] on natural language in virtual environment creation. Building on this, our work further explores fuzzy and more complex scenarios involving the repositioning of multiple objects.
V-A1 System Configuration
@@ V-B Benchmarking
V-B1 Baseline Systems
One of our baseline systems, for comparison is one of the publicly available prompts from NVIDIA’s Omniverse [^1] https://github.com/NVIDIA-Omniverse/kit-extension-sample-airoomgenerator/blob/main/exts/omni.example.airoomgenerator/omni/example/airoomgenerator/prompts.py. This prompt operates by generating a JSON output that details object placements within a 3D space, similar to our system. Contrary to Shenlong, the Omniverse prompt imposes no constraints on the reasoning scheme that the LLM should follow, requesting only the final positions of the objects. The Omniverse prompt accepts input specifications for each object, including its name, dimensions along the $X$, $Y$ and $Z$ axes, and a centrally located origin point. To ensure this baseline prompt is comparable to our system, we include only points of interest relevant to the scene editing query, rather than the entire scene description as initially done. This refined input approach significantly impacts the LLM’s response time. Our experiments demonstrate that providing only the relevant points of interest leads to a substantial decrease in response time—specifically, an average of 3.3±0.13.3$\pm$ 0.13.3 ± 0.1 times faster—without compromising overall accuracy. Due to these findings, we report results exclusively from this optimized methodology. Finally, we have extended the prompt by incorporating Euler angles for object orientation.
Figure 3: The left chart compares the average success rates, and the right chart compares the average LLM response times for three different systems—Omniverse, Euclidean, and CGA—across five categories of queries: Simple Queries, Compositional Queries, Fuzzy Queries, Compositional Fuzzy Queries, and Hard Queries. Success rates are measured on a scale from 0.0 to 1.0. Horizontal dashed lines indicate the overall mean success rate and response times for each system.
Our second baseline model is derived from recent work showing that LLMs can predict a dense sequence of end-effector poses for manipulation tasks [32]. We extend this concept to object repositioning on 3D scenes, treating it as a form of zero-shot trajectory generation. In this context, we task the LLM with constructing the necessary translation and rotation matrices to define the final trajectory, adapting the underlying techniques to suit scene manipulation objectives. We consider this baseline model to be the closest to our approach, as it guides the LLM to operate using templated functions that correspond to translation and rotation functions within the Euclidean space. This method aligns with our use of structured prompts that direct the LLM to generate specific geometric transformations necessary for scene editing tasks.
@@ V-B2 Performance Metrics
$S=\frac{1}{M}\sum_{i=1}^{M}\chi_{\text{correct}}(i)$ (9)
@@ V-B3 Results & Discussion
6. For coordinate extraction, employ the inner product operator |. For instance, use (X1.max | e2) * e2 to extract the Y-axis coordinate.
7. CONSIDER the current position of objects when moving one in relation to another. Calculate transformations using the initial locations derived from their bounding box points.
@@
- Its **IMPORTANT** to take into consideration all axis when performing relative transformations. Focus on the X **AND** Z axis when moving objects around the room. Make changes in the Y axis when a vertical movement is required. So, for example, when moving an object next to another, you should move it in both the X and Z axis, and when moving an object on top of another, you should move it in the X and Z axis, and also in the Y axis.
- When no axis is defined for rotations, the default axis is the Y-axis which means a rotation in the e1, e3 planes.
- When you want to move or rotate an object along an axis you must use the inner product operator |. For example, to move an object along the Y-axis you must use (X1.max | e2) * e2.
Figure 4: Template for Conformal Geometric Algebra Prompts used by Shenlong.
Performance Analysis by Query Group Figure 3 presents a detailed evaluation of the prompts’ performance across different groups of queries. Each group’s results are analyzed to highlight specific strengths and weaknesses of the system in handling varying complexities. Specifically, the effectiveness and efficiency of the CGA, Omniverse, and Euclidean prompts were evaluated across the five categories of queries: Simple, Compositional, Fuzzy, Compositional Fuzzy, and Hard. We focused on two key performance metrics: average success rate and average response time, providing insights into each system’s scene editing capabilities.
Overall, CGA achieves the highest average success rate of 0.80 and the lowest average response time of 18 seconds (p $<$ 0.05), with the results being statistically significant. Interestingly, all prompts exhibited the same overall standard deviation, both in terms of success rate and response time, indicating consistent variability across the different editors. To assess the differences between the systems, we performed Student’s t-tests. In comparison, the Omniverse and Euclidean prompts achieve success rates of 0.74 and 0.72, and average response times of 21.5 seconds and 24.5 seconds, respectively. While the differences in success rates among Omniverse and Euclidean prompts are not statistically significant, the variations in average response times are. We conjecture that the lack of statistical significance in terms of success rates among the systems can be attributed to the nature of the reasoning methods employed. Both the Omniverse and Euclidean prompts rely on classical geometrical reasoning, which may lead to similar levels of performance. In contrast, Shenlong utilizes a different approach to reasoning, which not only reflects in its superior performance but also in its statistical significance. This suggests that the distinct reasoning method employed by our system may contribute to its enhanced effectiveness.
@@
VI Limitations & Future Work
@@
Object Representations Currently, our system utilizes multiple intermediate representations of objects, including textual, templated, and bounding box references, requiring exact object names. To enhance this, we plan to implement semantic similarity measures at the token level or more advanced similarity searches using distributed representations. Additionally, we aim to enrich the system with further object information, such as orientation information, to better handle sophisticated queries involving relative rotations. By improving the system’s grasp of spatial relationships, improved performance in complex scenarios is expected.
@@
System Optimisations Since we preprocess and create templated queries, it is straightforward to perform query caching optimizations [36]. We aim to implement query caching optimisations to enhance both the accuracy and responsiveness of our system. We also envision transforming the collision module into a multimodal agent [51], enabling nuanced handling of complex scenarios. The agent could be provided with a top-down view of the scene to facilitate finding better resolutions. Lastly, we plan to integrate speech interaction capabilities into our system.